On January 1, 2015, Randy Incorporated purchased of $500,000 of 20 year, 10% bonds when the market rate of interest was 8% Interest is to be paid on June 30 and December 31 of each year.

1.Prepare the journal entry to record the purchase of the debt security classified as held to maturity.
2. Prepare the journal entry to record the receipt of the first two interest payments, assuming that Randy accounts for the debt security as held to maturity and uses the effective-interest method.
Answer #1

Solutions:

Journal Entries
Date Particulars Debit Credit
Jan 01 2015 Investment in Bonds $5,98,963.87
    Cash $5,98,963.87
To record purchase of 10% 500000 20 year bonds - Interest payable semi annually
June 30 2015 Cash 25000
Interest Income 23958.55
Investment in Bonds 1041.45
To record interest income { 598963*4% = 23958} and amortisation (25000-23958.55 = 1041.45)
Dec 31 2015 Cash 25000
Interest Income 23916.90
Investment in Bonds 1083.10
To record interest income { (598963-1041.45)*4% = 23916} and amortisation (25000-23916 = 1083.10)

Calculations

Present Value of Face Value of Bond $1,04,144.52
Present Value of Interest Payments $4,94,819.35
Total Amount Paid $5,98,963.87
Face Value of Bond $5,00,000.00
Initial Amount of Discount/(Premium) ($98,963.87)
Present Value of a Bond Maturity Value (Face Value)
Formula
Pvo FVn / (1+i)^n
FVn $5,00,000.00 Future Value
i 4% Semi Annual interest rate
n 40 Semi Annual Periods
(1+i) 1.04
(1+i)^n 4.8010
Pvo $1,04,144.52
Present Value of Semi Annual Interest Payment
Pmt 25000 Payments of a fixed amount
i 4.00% Interest Rate
n 40 no of payment period
Present Value of an Annuity
Formula
Present Value = PMT[1-1/(1+i)^n]/i
1+i 1.0400
(1+i)^n 4.8010
1/(1+i)^n 0.2083
1-1/(1+i)^n 0.7917
[1-1/(1+i)^n / i] 19.7928
Present Value $4,94,819.35